Systems Staff Meeting Notes 1/31/00

Systems Staff Meeting Notes
January 31, 2000

Present: Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Ying Tang
Announcements
  1. Bobbie's stepdaughter placed 9th in her weight class at a state weightlifting tournament over the weekend and her granddaughter has gotten her first tooth.
  2. Rat status: Ricky not caught in our traps yet. Bill found a suspiciously shredded piece of tinfoil in his office late Friday evening.
  3. Everett emailed Bobbie - he is enjoying his vacation & placed well in his races.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Two fiber farm switches on order, but not received.
  2. One work station for Facilities has been ordered.
  3. Eight work stations for DLC on order.
Liaisons
  1. Regular meeting scheduled this week (MMM as well).
Networking
  1. Two switches have been installed in Resource Services. The third floor should be finished today. The switches should eliminate the timeout problems some staff members were experiencing.
  2. NERDC reconfigured Journalism (which had the oldest 100 mg feeds on campus). They currently peak at 30 mg bandwidth, averaging under 256K. Journalism would like Gig-a-Bit feeds from the CORE. Our feeds from the CORE will be separated - put on a fractional subnet (16 or 32 address space). NERDC has proposed a B-POP, an active building POP device to be controlled by Net Services - rather than building administration. NERDC would have separate LAN subnets from ours, no cross-talk. They are also proposing feeds to the switches (from B-POPs) controlled by Building Domain. This has been proposed only for Journalism, hopefully not for us. Might consider it for Music, Education & AFA - but it would not be good for East, West or MSL.
  3. Network documentation - Lana Jamerson has been hired to work for Will, on network documentation. She should finish getting signed up this afternoon and be ready to start next week. Her schedule will be very flexible, not to run over 10 hours per week. She will be hired for the spring semester only.
  4. West wiring - meeting scheduled for tomorrow -Systems, West and Digital personnel will all attend. Data system, phone wiring and Com closets will all be added. There is a key issue with PPD regarding the Com closets that needs to be resolved. Wiring work should begin soon after the meeting. Once the phone wiring is in place, Patrick and his crew will need to do the phone installation. Bobbie mentioned that they could use phones in the Com closets. Bill will bring it up at the meeting.
  5. Proxy server hang-up this morning - by the time Winston & Bobbie checked on it, the problem resolved itself. It may be the throttle back at the CORE - we need to start maintaining time logs to see if there is an unusual burst of activity when the problem occurs. If we can get a precise time, then we can look at the network statistics. If our browser points to the proxy server & it works, then it can't be a proxy server problem. The TCP-IP interface on the mainframe has been problematic, but it shouldn't be affecting the proxy server. Could be a choke between the MSL network and the proxy server itself. Bill recently increased the filtered sites by two (braveworld & usa.com), but that should not affect the proxy server performance.
  6. The teleprocessing services at NERDC is now termed NS (Network Services).
Software Projects Status
  1. Q&A database replacement is a current project.
  2. OPS monthly reports on the Web - current.
  3. Personnel database needs change: need to be able to add one person with multiple jobs - there currently is no entry mechanism.
  4. CAT Desk Plus - ongoing dispute with the Library of Congress. It is working right now, except for printing.
Calendar
  1. Everett returns tomorrow, Feb. 1.
  2. Winston out next week, Feb. 7-11.
  3. Suzy out Mon-Wed, Feb. 7-9.
  4. Winston out Feb. 16-18.
  5. Hilton out Feb. 17.


Back to the Systems Homepage

Last updated February 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu